OpenStudy (anonymous):

solve 7y – 11 – 10y + 3 = 7 @KamiBug @nermaljean99 help! (please)

OpenStudy (anonymous):

I fan and medal!

OpenStudy (anonymous):

-3y-8=7 -3y=15 y=-5